OpenStudy (anonymous):

What two numbers have a sum = 5 and a product = -50 ?

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (compassionate):

Sum = when two numbers are added together Product = when two numbers are multiplied So sum = 5 product = -50 You need to find the factors of -50 that equal 5 when added.

OpenStudy (anonymous):

So like this: 5: 5 10 15 20 25 30... 50: 50 100 150 200...

OpenStudy (anonymous):

So 5 10 15 20 25 30 35 40 45 50 50 100 150 200

OpenStudy (anonymous):

But what do I do from here?

OpenStudy (compassionate):

No. So -50 factors are -25 x 2 (But no matter how we do this, when added, it will not equal 5) -5 * 10 = -50 Okay, and 10 - 5 (-5 + 10) = 5

OpenStudy (anonymous):

OK I get it. So the answer is -5 and 10

OpenStudy (compassionate):

Basically, the first one was to show you trial and error. While -25 x 2 does = -50, -25 + 2 does not equal 5. So second one is the corect one. You just gotta think about these types of things. (: